# Trellis codes with a good distance profile constructed from expander graphs

### v1

@@ I-B Trellis codes
The rate of an $(n,M)$ trellis code over $\Sigma_{q}$ is defined to be $\log_{q}M/n$. Note that a trellis code may admit multiple generator digraphs, and thus the number of states needed to describe the trellis code need not be unique. With the operational meaning of the degree of convolutional codes in mind, we define the degree of a trellis code $\mathscr{C}$} over $\Sigma_{q}$ as $\delta({\mathscr{C}})=\min\{\log_{q}|V|:\Gamma=(V,E,\psi)\text{ is a generator digraph of }{\mathscr{C}}\}$.
@@
$d({\mathscr{C}})=\min_{{c,c^{\prime}\in{\mathscr{C}},c\neq c^{\prime}}}d(c,c^{\prime}).$
@@ III-B Code parameters
${d_{j}^{c}(\tilde{{\mathscr{C}}}_{\phi})}={\sum_{i=0}^{j}|S_{i}|}.$
@@
$\frac{d_{j}^{c}(\tilde{{\mathscr{C}}}_{\phi})}{(j+1)n}$ $\geq\frac{\sum_{i=0}^{j}\operatorname{arw}(\bar{T}_{i})-(j+1)\gamma\sqrt{\theta^{-1}}}{(j+1)(1-\gamma)}.$
@@
The dimension of $\tilde{{\mathscr{C}}}$ is at least $(k-(m+1)(1-r)\Delta)n$ and the rate of $\tilde{{\mathscr{C}}}_{\phi}$ satisfies
@@
We begin by examining the dimension of $\mathscr{B}$}, which is equal to the dimension of $\tilde{{\mathscr{C}}}$. Note that ${\mathscr{B}}={\mathscr{B}}^{(1)}\cap{\mathscr{B}}^{(2)}$. Moreover, from (7) and (8), we have $\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}^{(1)}=n\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}_{1}=nk$ and $\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}^{(2)}=n\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}_{2}^{m+1}=n(m+1)r\Delta$. Therefore, the ${\mathbb{F}}_{q}$-dimension of $\mathscr{B}$} satisfies
@@
$\frac{\log_{q^{r\Delta}}|\tilde{{\mathscr{C}}}_{\phi}|}{n}=\frac{\log_{q^{r\Delta}}|\tilde{{\mathscr{C}}}|}{n}=\frac{\dim_{{\mathbb{F}}_{q}}\tilde{{\mathscr{C}}}}{nr\Delta}=\frac{\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}}{nr\Delta}\geq\frac{k}{r\Delta}-(m+1)\Big(\frac{1}{r}-1\Big).$
@@ Appendix D Proof of Lemma 9
where $T_{0}\cup\dots\cup T_{\ell-1}:=\emptyset$ if $\ell=0$. The above identity follows from noticing that one can partition the set $T_{0}\cup\dots\cup T_{j}$ into subsets $T_{0},T_{1}\setminus T_{0},\ldots,T_{j}\setminus(T_{0}\cup\dots\cup T_{j-1})$ and sum over each subset. For $\ell\in\llbracket{j}\rrbracket$, it holds that

### v2

@@ I-B Trellis codes
The rate of an $(n,M)$ trellis code $\mathscr{C}$} over $\Sigma_{q}$ is defined to be $\log_{q}M/n$. We refer to $M$ as the size of $\mathscr{C}$}, and write $|{\mathscr{C}}|:=M$. [^1] Similarly, for an $(n,k)$ convolutional code $\mathscr{C}$} over ${\mathbb{F}}_{q}$, we refer to $q^{k}$ as the size of $\mathscr{C}$} and write $|{\mathscr{C}}|=q^{k}$. Note that a trellis code may admit multiple generator digraphs, and thus the number of states needed to describe the trellis code need not be unique. With the operational meaning of the degree of convolutional codes in mind, we define the degree of a trellis code $\mathscr{C}$} over $\Sigma_{q}$ as $\delta({\mathscr{C}})=\min\{\log_{q}|V|:\Gamma=(V,E,\psi)\text{ is a generator digraph of }{\mathscr{C}}\}$.
@@
$d_{\mathrm{free}}({\mathscr{C}})=\min_{{c,c^{\prime}\in{\mathscr{C}},c\neq c^{\prime}}}d(c,c^{\prime}).$
@@ III-B Code parameters
${d_{j}^{c}(\tilde{{\mathscr{C}}}_{\phi})}={\min_{c(D)\in\tilde{{\mathscr{C}}}:c_{0}\neq 0}}{\sum_{i=0}^{j}|S_{i}|}.$
@@
$\frac{d_{j}^{c}(\tilde{{\mathscr{C}}}_{\phi})}{(j+1)n}$ $\geq{\min_{c(D)\in\tilde{{\mathscr{C}}}:c_{0}\neq 0}}\frac{\sum_{i=0}^{j}\operatorname{arw}(\bar{T}_{i})-(j+1)\gamma\sqrt{\theta^{-1}}}{(j+1)(1-\gamma)}.$
@@
The ${\mathbb{F}}_{q}[D]$-rank [^2] This ${\mathbb{F}}_{q}[D]$-rank is often informally referred to as the ${\mathbb{F}}_{q}$-dimension of convolutional codes over ${\mathbb{F}}_{q}$. of $\tilde{{\mathscr{C}}}$ is at least $(k-(m+1)(1-r)\Delta)n$ and the rate of $\tilde{{\mathscr{C}}}_{\phi}$ satisfies
@@
We begin by examining the ${\mathbb{F}}_{q}$-dimension of $\mathscr{B}$}, which is equal to the ${\mathbb{F}}_{q}[D]$-rank of $\tilde{{\mathscr{C}}}$. Note that ${\mathscr{B}}={\mathscr{B}}^{(1)}\cap{\mathscr{B}}^{(2)}$. Moreover, from (7) and (8), we have $\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}^{(1)}=n\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}_{1}=nk$ and $\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}^{(2)}=n\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}_{2}^{m+1}=n(m+1)r\Delta$. Therefore, the ${\mathbb{F}}_{q}$-dimension of $\mathscr{B}$} satisfies
@@
$\frac{\log_{q^{r\Delta}}|\tilde{{\mathscr{C}}}_{\phi}|}{n}=\frac{\log_{q^{r\Delta}}|\tilde{{\mathscr{C}}}|}{n}=\frac{{\operatorname{rank}_{{\mathbb{F}}_{q}[D]}}\tilde{{\mathscr{C}}}}{nr\Delta}=\frac{\dim_{{\mathbb{F}}_{q}}{\mathscr{B}}}{nr\Delta}\geq\frac{k}{r\Delta}-(m+1)\Big(\frac{1}{r}-1\Big).$
@@ Appendix D Proof of Lemma 9
where $T_{0}\cup\dots\cup T_{\ell-1}:=\emptyset$ if $\ell=0$. The above identity follows from noticing that one can partition the set $T_{0}\cup\dots\cup T_{j}$ into subsets $T_{0},T_{1}\setminus T_{0},\ldots,T_{j}\setminus(T_{0}\cup\dots\cup T_{j-1})$ and sum over each subset. [^4] We adopt the convention that a sum over the empty set is zero. For $\ell\in\llbracket{j}\rrbracket$, it holds that